In the capacitor arrangement shown C1 = 75 mF andIn the capacitor arrangement shown C1 = 75 mF and C2 = 41 mF. If a voltage of V = 98 volts is applied, what is the energy stored (J) in the capacitor C1?
Answer #1

Voltage across C1 is

V1 =V*(C2/C1+C2)

V1=98*(41/41+75)

V1=34.64 Volts

Energy stored in C1 is

E1=(1/2)C1*V1^2 =(1/2)*(75*10^-3)*34.64^2

E1=44.99 J=45 J(approx)